OpenStudy (anonymous):

Given: x + y = 10. If x = -21, what is y?

Vocaloid (vocaloid):

replace "x" with "-21" -21 + y = 10 y = ?

OpenStudy (anonymous):

31

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@Vocaloid how do i find what y is?

Vocaloid (vocaloid):

-21 + y = 10 start by adding 21 to both sides

OpenStudy (anonymous):

10+21=31

OpenStudy (anonymous):

i got 31 @Vocaloid

OpenStudy (anonymous):

thats correct

Vocaloid (vocaloid):

good!

OpenStudy (anonymous):

@Vocaloid thank you!